OpenStudy (anonymous):

Call Wireless phone company charges $0.50 for the first call you make and $0.10 for each additional minute. The cost of a phone call can be expressed by the formula C = 0.50 + 0.10(m – 1), where C is the total cost in dollars and m is the number of minutes that you talk. If the bill for your phone call was $1.90, how long did you talk? A. 12 minutes B. 14 minutes C. 15 minutes D. 16 minutes

OpenStudy (anonymous):

help

OpenStudy (anonymous):

B

OpenStudy (anonymous):

1.90 = 0.50 + 0.10(m-1) 1.40 = 0.10(m-1) 14.0 = m-1 m = 14 + 1 = 15 check: 0.50 + 0.1(14)= 1.90 yes